OpenStudy (anonymous):

Find equations of both the tangent lines to the ellipse x2 + 9y2 = 81 that pass through the point (27, 3).

OpenStudy (anonymous):

x^2/81 + y^2/9 = 1. so the tangent is 27x/81 +3y/9 = 1 or x + y - 3 = 0

OpenStudy (anonymous):